OpenStudy (anonymous):

HELP MEDAL GOES TO PERSON THAT GETS IT RIGHT!!!!! The length of the diagonal of a square is 8 inches. What is the area of the square?

OpenStudy (magbak):

The area is 4 square inches.

OpenStudy (anonymous):

4

OpenStudy (magbak):

You copied me.

thomaster (thomaster):

Don't just give the answer. guide her to the answer...

OpenStudy (magbak):

I got it first.

OpenStudy (anonymous):

4 is not an answer choice

OpenStudy (magbak):

Thank you is none of the above an answer choice.

OpenStudy (anonymous):

none of the above is not an answer choice. either

OpenStudy (anonymous):

2

OpenStudy (magbak):

What are the answer choices

OpenStudy (anonymous):

The length of the diagonal of a square is 8 inches. What is the area of the square? Question 11 options: 16 square root(2) sq. in. 32 sq. in. 64 square root(2) sq. in. 128 sq. in.

thomaster (thomaster):

Area of a square is half the product of the diagonals. So 8*8 = 64 64/2 = 32 square inches

OpenStudy (anonymous):

2*8 =16

OpenStudy (magbak):

IT is a

OpenStudy (magbak):

It is the same as 4 exactly.

OpenStudy (magbak):

jim_thompson5910

OpenStudy (magbak):

@ jim_thompson5910

OpenStudy (magbak):

@jim_thompson5910